OpenStudy (anonymous):

Each of the numbers from 1 through 30 is written on a table tennis ball and placed in a wire cage. Each of the numbers 20 through 45 is written on a table tennis ball and placed in a different wire cage. One ball is chosen at random from each spinning cage. Which of the following best describes: P(each is a 25)

OpenStudy (kropot72):

The probability of drawing ball number 25 from the cage with balls numbered 1 through 30 is 1/30. The probability of drawing ball number 25 from the cage with balls numbered 20 through 45 is 1/26. \[P(each\ is\ a\ 25=\frac{1}{30}\times\frac{1}{26}\]